OpenStudy (anonymous):

Solve this equation 5^ (x-2) = 3^(3x+2) I did on the calculator and got -3.211636. Is there a more sensible answer?

OpenStudy (amistre64):

im thinking if i see it right: (x-2) log5(5) = (3x+2) log5(3) (x-2)/log5(3) = 3x+2 x/log5(3) - 2/log5(3) = 3x+2 x/log5(3) - 3x = 2 + 2/log5(3) x(1/log5(3) - 3) = 2 + 2/log5(3) x = (2 + 2/log5(3))/(1/log5(3) - 3) maybe? lol

OpenStudy (amistre64):

-3.2116....

OpenStudy (anonymous):

the eff